Could you please tell me how to set up the correct equation for this word problem. "you are 13 years older than your sister, in nine years you will be twice as old as your sister. How old is your sister".
I know that 13+9=22, if i divide that by 2 then she would be 11. I do not understand how to put this into a equation in order to solve it.

Answer by htmentor(1343)

Let S = your sister's current age
Then your current age = S + 13
In 9 years, your age will be S + 13 + 9 and your sister's age will be S + 9
Given: In 9 years your age will be twice that of your sisters, so
S + 22 = 2(S + 9)
Solve for S:
S = 22 - 18 = 4
So your sister is 4 years old